# GRVFL-2V: Graph Random Vector Functional Link Based on Two-View Learning

### v1

@@ 3.1 LFDA Under the GE Framework
Algorithm 1 GRVFL-2V Model Algorithm
@@
Output: GRVFL-2V model.
@@
4 Computational Complexity
We analyze the computational complexity of our proposed GRVFL-2V model in this section. For $X_{A}$, computing the graph embedding (GE) matrix $G_{1}$ using the methodologies from [52] which account for intrinsic as well as penalty graph structures results in a time complexity of $\mathcal{O}((m+h_{1})^{3}+(m+h_{1})^{2}l)$. Similarly, for $X_{B}$, computing $G_{2}$ yields a complexity of $\mathcal{O}((n+h_{2})^{3}+(n+h_{2})^{2}l)$. Thus, the GE process incurs a total computational complexity of $\mathcal{O}((n+h_{1})^{3}+(n+h_{1})^{2}l)+\mathcal{O}((m+h_{2})^{3}+(m+h_{2})^{2}l)$. To solve our model, we address (25). The complexity of this step is primarily governed by the inversion of a square matrix with an order of $(m+n+h_{1}+h_{2})$, leading to a computational complexity of $\mathcal{O}((m+n+h_{1}+h_{2})^{3})$. Consequently, the total computational complexity of our proposed model becomes $\mathcal{O}((n+h_{1})^{3}+(n+h_{1})^{2}l)+\mathcal{O}((m+h_{2})^{3}+(m+h_{2})^{2}l)+\mathcal{O}((m+n+h_{1}+h_{2})^{3})\approx\mathcal{O}((m+n+h_{1}+h_{2})^{3})$.
@@ 5 Experiments and Result Discussion
The performance of the proposed model has been evaluated against the baseline models: SVM2K [41], MvTSVM [42], ELM1 (performance of ELM [13] on ‘$view-A$’), ELM2 (performance of ELM [13] on ‘$view-B$’), RVFL1 (performance of RVFL [11] on ‘$view-A$’) and RVFL2 (performance of RVFL [11] on ‘$view-B$’), and MVLDM [53]. The datasets used for the evaluation are UCI [54], KEEL [55], Animal with Attributes (AwA) [^1] http://attributes.kyb.tuebingen.mpg.de, and Corel5K [^2] https://wang.ist.psu.edu/docs/related/.
@@ 5.2 Experiments on UCI and KEEL Datasets
Table 1: Average Accuracy and Average Rank of the baseline models and the proposed GRVFL-2V model over UCI and KEEL datasets.
Model SVM2K [41] MvTSVM [42] ELM1 [13] ELM2 [13] RVFL1 [11] RVFL2 [11] MVLDM [53] GRVFL-2V (proposed)
Average ACC 75.24 65.83 82.11 80.94 82.63 80.26 80.58 84.35
Average Rank 5.52 7.69 4.02 4.30 3.54 4.41 4.80 1.74
@@
Within this subsection, we delve into analyzing the statistical significance of the results obtained from our experiments, specifically concentrating on datasets sourced from UCI and KEEL repositories [54, 55]. Through our assessment, we encompass a total of 27 datasets. Given the absence of inherent multi-view characteristics in the UCI and KEEL datasets, we designated the 95% principal component extracted from the original data as ‘$view-B$’, while the unaltered data itself acts as ‘$view-A$’. In order to thoroughly compare the performance levels between our proposed graph random vector functional link based on the two-view learning (GRVFL-2V) model and the baseline models with optimized hyperparameters, please refer to Table S.I in the Supplementary Material. The outcomes illustrated in Table 1 in the paper underline the average accuracy (ACC) and average rank of each model across the UCI and KEEL datasets. The baseline models SVM2K, MvTSVM, ELM1, EML2, RVFL1, RVFL2, and MVLDM achieves the average ACC of $75,24\%$, $65.83\%$, $82.11\%$, $80.94\%$, $82.62\%$, $80.26\%$, and $80.58\%$, respectively. The GRVFL-2V model that we put forward exhibited an outstanding average accuracy rate of $84.35\%$, surpassing the performance levels of the baseline models. The difference in average ACC of the proposed model and the baseline models lies approximately between $4\%-20\%$. This showcases the superior generalization capabilities inherent in our proposed model compared to the baseline models.
@@
In order to further assess the effectiveness of the proposed model, a ranking method is utilized. This method involves assigning a rank to each model for every dataset, where the model that performs the best is given the lowest rank, and the model that performs the worst is given the highest rank. The average rank for each model is then calculated by finding the mean of its ranks across all datasets. If there are a total of $N$ datasets, each evaluated with $\lambda$ models, the rank of the $p^{th}$ model on the $t^{th}$ dataset can be represented as ${s^{t}_{p}}$. Subsequently, the average rank of the $p^{th}$ model is calculated as ${R}_{p}=\frac{1}{N}\sum_{t=1}^{N}s^{t}_{p}$. The proposed GRVFL-2V model has been able to achieve an average rank of 1.74. On the other hand, the baseline models such as SVM2K, MvTSVM, ELM1, ELM2, RVFL1, RVFL2, and MVLDM have average ranks of 5.52, 7.69, 4.02, 4.30, 3.54, 4.41, and 4.80, respectively. These results clearly demonstrate the superiority of the proposed model over the baseline models.
@@
In order to assess the statistical significance of the proposed model, the Friedman test is utilized as outlined in [56]. This particular test is designed to pinpoint noteworthy variations among the models being compared by scrutinizing their average ranks. The underlying assumption of the null hypothesis is that all models showcase an identical average rank, indicating an equivalent level of performance. The Friedman test adheres to a chi-squared distribution denoted as $\chi^{2}_{F}$ with $(\lambda-1)$ degrees of freedom, which can be calculated using the formula: $\chi^{2}_{F}=\frac{12N}{\lambda(\lambda+1)}\left[\sum_{p}{{R}_{p}}^{2}-\frac{\lambda(\lambda+1)^{2}}{4}\right]$. Here, ${R}_{p}$ signifies the average rank of the $p$th model, $\lambda$ represents the total number of models, and $N$ denotes the number of datasets involved in the analysis. The Friedman statistic $F_{F}$ is determined by the formula: $F_{F}=\frac{(N-1)\chi^{2}_{F}}{N(\lambda-1)-\chi^{2}_{F}}$, where the $F$-distribution is characterized by $(\lambda-1)$ and $(\lambda-1)\times(N-1)$ degrees of freedom. Given that $N=27$ and $\lambda=8$, the values of $\chi^{2}_{F}$ and $F_{F}$ are calculated to be 91.37 and 24.33, respectively. By consulting the $F$-distribution table at a 5% level of significance $(\alpha)$, the critical value of $F_{F}(7,182)=2.39$ is obtained. Since the calculated value of $F_{F}$ is 24.33, which exceeds the critical value of 2.39, the null hypothesis is rejected. This outcome underscores a substantial statistical distinction among the models under comparison.
@@
Table 2: The statistical comparison of the proposed GRVFL-2V model with the baseline models on UCI and KEEL datasets using the Nemenyi post hoc test.
SVM2K [41] MvTSVM [42] ELM1 [13] ELM2 [13] RVFL1 [11] RVFL2 [11] MVLDM [53]
GRVFL-2V (proposed)
@@
✓  indicates that the model listed in the row is superior to the model mentioned in the column.
@@
In addition, the Nemenyi post hoc test [56] is utilized to delve deeper into the differences between the models. If the average ranks of the models exhibit a variance of at least the critical difference ($CD$), then they are deemed to be significantly distinct. The $CD$ is determined by the formula: $CD=q_{\alpha}{\begin{bmatrix}\frac{\lambda(\lambda+1)}{6N}\end{bmatrix}}^{1/2}$. On calculating, we get the $CD=1.35$. The results presented in Table 2 provide clear evidence that the average rank difference between the baseline model and the proposed GRFVL-MVL is larger than the critical difference ($CD$) value. To be more precise, the average rank difference between MvTSVM and GRVFL-2V is $5.95$, which is greater than the $CD$ value of $1.35$. Similarly, the average rank difference between MVLDM and GRVFL-2V is $3.06$, also surpassing the $CD$ value. Therefore, these results unequivocally demonstrate that the proposed model exhibits significant dissimilarity compared to the baseline models.
@@
Furthermore, a pairwise win-tie-loss sign test is conducted under the assumption of equal performance between the two models under the null hypothesis. It is anticipated that each model will emerge victorious in roughly half of the total datasets, denoted as $\frac{N}{2}$, where $N$ represents the total number of datasets. To establish statistical significance, a model must secure wins in approximately $\frac{N}{2}+1.96\frac{\sqrt{N}}{2}$ datasets more than its counterpart. In scenarios where there is an even number of ties between the models, the ties are evenly distributed. Conversely, in cases of an odd number of ties, one tie is discounted, and the remaining ties are divided between the models. With $N$ set at 27, a minimum of 18.59 wins is required to establish a significant difference between the two models.
@@
Table 3: Pairwise win-tie-loss test of proposed GRVFL-2V model and baseline models on UCI and KEEL datasets.
SVM2K [41] MvTSVM [42] ELM1 [13] ELM2 [13] RVFL1    [11] RVFL2   [11] Large_Margin    [53]
@@
The results depicted in Table 3 showcase a distinct advantage for our proposed model over the baseline model. Clearly, we can see that our proposed model wins on at least 21 ($>$ 18.59) datasets out of 27 datasets. Hence, our proposed model is far superior to the baseline models. In particular, we can observe that our proposed GRVFL-2V surpasses SVM2K in 25 datasets out of the 27 datasets. Furthermore, it outperforms the MvTSVM in all 27 datasets while excelling against MVLDM in 25 datasets out of the 27. The findings unequivocally affirm that our proposed model exhibits a significantly higher performance level compared to the baseline models.
@@ 5.3 Experiments on Corel5k Datasets
Table 4: Average Accuracy and Average Rank of the baseline models and the proposed model GRVFL-2V over Corel5k datasets.
Model SVM2K [41] MvTSVM [42] ELM1 [13] ELM2 [13] RVFL1 [11] RVFL2 [11] MVLDM [53] GRVFL-2V (proposed)
@@
The results presented in Table 4 provide clear evidence that our proposed model outperforms the baseline models. With an average accuracy (ACC) of 77.33%, our model achieves the highest accuracy among all the models compared. Additionally, the proposed GRVFL-2V model obtains the lowest average rank of 2.94, indicating its superior performance compared to the baseline models. These findings are from the 50 datasets from Corel5k, which further solidifies the superiority of our proposed model. For more detailed experimental results, please refer to Table S.II in the Supplementary Material.
@@ 5.4 Experiments on Animal with Attributes (AwA) Datasets
Table 5: Average Accuracy and Average Rank of the baseline models and the proposed model GRVFL-2V over AwA datasets.
Model SVM2K [41] MvTSVM [42] ELM1 [13] ELM2 [13] RVFL1 [11] RVFL2 [11] MVLDM [53] GRVFL-2V (proposed)
@@
The average rank and ACC of our proposed model are displayed in Table 5 in relation to the baseline models. Across the 45 datasets from AwA, our model achieved the highest average accuracy and the lowest average rank. The average ACC of proposed GRVFL-2V is 83.29, which is upto 4% to 20% higher than the baseline models. Also, GRVFL-2V has the lowest average rank of 1.56. The results clearly demonstrate the superior performance of the proposed GRVFL-2V model over the baseline models. For more detailed experimental findings, please refer to Table S.III in the Supplementary Material.
@@ 5.5 Sensitivity Analysis of Hyperparameters $C_{1}$ and $C_{2}$
Figure 2: The effect of hyperparameter $(c_{1},c_{2})$ tuning on the accuracy (ACC) of some UCI and KEEL datasets on the performance of proposed GRVFL-2V model.
To gain a comprehensive understanding of how hyperparameters affect the generalization ability of the proposed GRVFL-2V model, we conducted a systematic exploration of the hyperparameter space by tuning the values of $c_{1}$ and $c_{2}$. This helps us to identify the optimal configuration that maximizes predictive accuracy and improves the model’s resilience to previously unseen data. Fig. 2. visually represents how the accuracy of the model behaves when the hyperparameters are tuned. The visual clearly shows that the proposed model is highly sensitive to the values of hyperparameters $c_{1}$ and $c_{2}$. In Fig. 2. (a), we see that optimal accuracy is achieved when $c_{1}=10^{5}$ and $c_{2}=10^{4}$ whereas, in Fig. 2. (c), optimal configuration is found at two distinct coordinates, $(10^{5},10^{5})$ and $(10,10^{5})$. Similar observations can be made for Fig. 2. (b) and Fig. 2. (d). Overall, these findings underscore the need for careful selection of hyperparameter values to achieve optimal model performance.

### v3

@@ 3.1 LFDA Under the GE Framework
Algorithm 1 GRVFL-MV Model Algorithm
@@
Output: GRVFL-MV model.
@@
3.2 Computational Complexity
We analyze the computational complexity of our proposed GRVFL-MV model in this section. For $X_{A}$, computing the graph embedding (GE) matrix $G_{1}$ using the methodologies from [16] which account for intrinsic as well as penalty graph structures results in a time complexity of $\mathcal{O}((m+h_{1})^{3}+(m+h_{1})^{2}l)$. Similarly, for $X_{B}$, computing $G_{2}$ yields a complexity of $\mathcal{O}((n+h_{2})^{3}+(n+h_{2})^{2}l)$. Thus, the GE process incurs a total computational complexity of $\mathcal{O}((n+h_{1})^{3}+(n+h_{1})^{2}l)+\mathcal{O}((m+h_{2})^{3}+(m+h_{2})^{2}l)$. To solve our model, we address (25). The complexity of this step is primarily governed by the inversion of a square matrix with an order of $(m+n+h_{1}+h_{2})$, leading to a computational complexity of $\mathcal{O}((m+n+h_{1}+h_{2})^{3})$. Consequently, the total computational complexity of our proposed model becomes $\mathcal{O}((n+h_{1})^{3}+(n+h_{1})^{2}l)+\mathcal{O}((m+h_{2})^{3}+(m+h_{2})^{2}l)+\mathcal{O}((m+n+h_{1}+h_{2})^{3})\approx\mathcal{O}((m+n+h_{1}+h_{2})^{3})$.
@@ 4 Experiments and Result Discussion
The performance of the proposed model has been evaluated against the baseline models: SVM2K [8], MvTSVM [39], RVFLwoDL1 (RVFL without direct link [14] on ‘$view-A$’), RVFLwoDL2 (RVFL without direct link [14] on ‘$view-B$’), RVFL1 (RVFL [24] on ‘$view-A$’) and RVFL2 (RVFL [24] on ‘$view-B$’), and MVLDM [13]. Theoretical comparisons of the proposed model against the baseline models are made in the supplementary Section S.I. The datasets used for the evaluation are UCI [7], KEEL [6], Animal with Attributes (AwA) [^1] http://attributes.kyb.tuebingen.mpg.de, and Corel5K [^2] https://wang.ist.psu.edu/docs/related/.
@@ 4.2 Experiments on UCI and KEEL Datasets
Table 1: Average accuracy and average rank of the baseline models and the proposed GRVFL-MV model over UCI and KEEL datasets.
Model SVM2K [8] MvTSVM [39] RVFLwoDL1 [14] RVFLwoDL2 [14] RVFL1 [24] RVFL2 [24] MVLDM [13] GRVFL-MV (proposed)
Average ACC $76.70$ $64.73$ $82.99$ $82.35$ $83.59$ $81.59$ $79.88$ $\mathbf{85.68}$
Average Rank $5.59$ $7.71$ $4.02$ $4.21$ $3.45$ $4.38$ $4.91$ $\mathbf{1.74}$
@@
Within this subsection, we delve into analyzing the statistical significance of the results obtained from our experiments, specifically concentrating on datasets sourced from UCI and KEEL repositories [7, 6]. Through our assessment, we encompass a total of 29 datasets. Given the absence of inherent multi-view characteristics in the UCI and KEEL datasets, we designated the 95% principal component extracted from the original data as ‘$view-B$’, while the unaltered data itself acts as ‘$view-A$’. In order to thoroughly compare the performance levels between our proposed graph random vector functional link based on the multi-view learning (GRVFL-MV) model and the baseline models with optimized hyperparameters, please refer to Table S.1 in the supplementary file. The outcomes illustrated in Table 1 in the paper underline the average accuracy (ACC) and average rank of each model across the UCI and KEEL datasets. The baseline models SVM2K, MvTSVM, RVFLwoDL1, RVFLwoDL2, RVFL1, RVFL2, and MVLDM achieves the average ACC of $76.70\%$, $64.73\%$, $82.99\%$, $82.35\%$, $83.59\%$, $81.59\%$, and $79.88\%$, respectively. The GRVFL-MV model that we put forward exhibited an outstanding average accuracy rate of $85.68\%$, surpassing the performance levels of the baseline models. The difference in average ACC of the proposed model and the baseline models lies approximately between $4\%-20\%$. This showcases the superior generalization capabilities inherent in our proposed model compared to the baseline models.
@@
In order to further assess the effectiveness of the proposed model, a ranking method is utilized. This method involves assigning a rank to each model for every dataset, where the model that performs the best is given the lowest rank, and the model that performs the worst is given the highest rank. The average rank for each model is then calculated by finding the mean of its ranks across all datasets. If there are a total of $N$ datasets, each evaluated with $\lambda$ models, the rank of the $p^{th}$ model on the $t^{th}$ dataset can be represented as ${s^{t}_{p}}$. Subsequently, the average rank of the $p^{th}$ model is calculated as ${R}_{p}=\frac{1}{N}\sum_{t=1}^{N}s^{t}_{p}$. The proposed GRVFL-MV model has been able to achieve an average rank of 1.74. On the other hand, the baseline models such as SVM2K, MvTSVM, RVFLwoDL1, RVFLwoDL2, RVFL1, RVFL2, and MVLDM have average ranks of 5.59, 7.71, 4.02, 4.21, 3.45, 4.38, and 4.91, respectively. These results clearly demonstrate the superiority of the proposed model over the baseline models.
@@
In order to assess the statistical significance of the proposed model, the Friedman test is utilized as outlined in [5]. This particular test is designed to pinpoint noteworthy variations among the models being compared by scrutinizing their average ranks. The underlying assumption of the null hypothesis is that all models showcase an identical average rank, indicating an equivalent level of performance. The Friedman test adheres to a chi-squared distribution denoted as $\chi^{2}_{F}$ with $(\lambda-1)$ degrees of freedom, which can be calculated using the formula: $\chi^{2}_{F}=\frac{12N}{\lambda(\lambda+1)}\left[\sum_{p}{{R}_{p}}^{2}-\frac{\lambda(\lambda+1)^{2}}{4}\right]$. Here, ${R}_{p}$ signifies the average rank of the $p$th model, $\lambda$ represents the total number of models, and $N$ denotes the number of datasets involved in the analysis. The Friedman statistic $F_{F}$ is determined by the formula: $F_{F}=\frac{(N-1)\chi^{2}_{F}}{N(\lambda-1)-\chi^{2}_{F}}$, where the $F$-distribution is characterized by $(\lambda-1)$ and $(\lambda-1)\times(N-1)$ degrees of freedom. Given that $N=29$ and $\lambda=8$, the values of $\chi^{2}_{F}$ and $F_{F}$ are calculated to be 100.5 and 27.45, respectively. By consulting the $F$-distribution table at a 5% level of significance $(\alpha)$, the critical value of $F_{F}(7,196)=2.05$ is obtained. Since the calculated value of $F_{F}$ is 27.45, which exceeds the critical value of 2.05, the null hypothesis is rejected. This outcome underscores a substantial statistical distinction among the models under comparison.
@@
Table 2: The statistical comparison of the proposed GRVFL-MV model with the baseline models on UCI and KEEL datasets using the Nemenyi post hoc test.
SVM2K [8] MvTSVM [39] RVFLwoDL1 [14] RVFLwoDL2 [14] RVFL1 [24] RVFL2 [24] MVLDM [13]
GRVFL-MV (proposed)
@@
✓  indicates that the model listed in the row is superior to the model mentioned in the column, while ✗ indicates otherwise.
@@
In addition, the Nemenyi post hoc test [5] is utilized to delve deeper into the differences between the models. If the average ranks of the models exhibit a variance of at least the critical difference ($CD$), then they are deemed to be significantly distinct. The $CD$ is determined by the formula: $CD=q_{\alpha}{\begin{bmatrix}\frac{\lambda(\lambda+1)}{6N}\end{bmatrix}}^{1/2}$. On calculating, we get the $CD=1.94$. The results presented in Table 2 provide clear evidence that the average rank difference between the baseline model and the proposed GRFVL-MVL is larger than the critical difference ($CD$) value. To be more precise, the average rank difference between MvTSVM and GRVFL-MV is $5.97$, which is greater than the $CD$ value of $1.94$. Similarly, the average rank difference between MVLDM and GRVFL-MV is $3.17$, also surpassing the $CD$ value. Therefore, these results unequivocally demonstrate that the proposed model exhibits significant dissimilarity compared to the baseline models.
@@
Furthermore, a pairwise win-tie-loss sign test is conducted under the assumption of equal performance between the two models under the null hypothesis. It is anticipated that each model will emerge victorious in roughly half of the total datasets, denoted as $\frac{N}{2}$, where $N$ represents the total number of datasets. To establish statistical significance, a model must secure wins in approximately $\frac{N}{2}+1.96\frac{\sqrt{N}}{2}$ datasets more than its counterpart. In scenarios where there is an even number of ties between the models, the ties are evenly distributed. Conversely, in cases of an odd number of ties, one tie is discounted, and the remaining ties are divided between the models. With $N$ set at 29, a minimum of 19.78 wins is required to establish a significant difference between the two models.
@@
Table 3: Pairwise win-tie-loss test of proposed GRVFL-MV model and baseline models on UCI and KEEL datasets.
SVM2K [8] MvTSVM [39] RVFLwoDL1 [14] RVFLwoDL2 [14] RVFL1    [24] RVFL2    [24] Large_Margin    [13]
@@
The results depicted in Table 3 showcase a distinct advantage for our proposed model over the baseline model. Clearly, we can see that our proposed model wins on at least 22 ($>$ 19.78) datasets out of 27 datasets. Hence, our proposed model is far superior to the baseline models. In particular, we can observe that our proposed GRVFL-MV surpasses SVM2K in 27 datasets out of the 29 datasets. Furthermore, it outperforms the MvTSVM in all 29 datasets while excelling against MVLDM in 27 datasets out of the 29. The findings unequivocally affirm that our proposed model exhibits a significantly higher performance level compared to the baseline models.
@@ 4.3 Experiments on Corel5k Datasets
Table 4: Average accuracy and average rank of the baseline models and the proposed model GRVFL-MV over Corel5k datasets.
Model SVM2K [8] MvTSVM [39] RVFLwoDL1 [14] RVFLwoDL2 [14] RVFL1 [24] RVFL2 [24] MVLDM [13] GRVFL-MV (proposed)
@@
The results presented in Table 4 provide clear evidence that our proposed model outperforms the baseline models. With an average accuracy (ACC) of 77.33%, our model achieves the highest accuracy among all the models compared. Additionally, the proposed GRVFL-MV model obtains the lowest average rank of 2.94, indicating its superior performance compared to the baseline models. These findings are from the 50 datasets from Corel5k, which further solidifies the superiority of our proposed model. For more detailed experimental results, please refer to Table S.2 in the supplementary file.
@@ 4.4 Experiments on Animal with Attributes (AwA) Datasets
Table 5: Average accuracy and average rank of the baseline models and the proposed model GRVFL-MV over AwA datasets.
Model SVM2K [8] MvTSVM [39] RVFLwoDL1 [14] RVFLwoDL2 [14] RVFL1 [24] RVFL2 [24] MVLDM [13] GRVFL-MV (proposed)
@@
The average rank and ACC of our proposed model are displayed in Table 5 in relation to the baseline models. Across the 45 datasets from AwA, our model achieved the highest average accuracy and the lowest average rank. The average ACC of proposed GRVFL-MV is 83.29, which is upto 4% to 20% higher than the baseline models. Also, GRVFL-MV has the lowest average rank of 1.56. The results clearly demonstrate the superior performance of the proposed GRVFL-MV model over the baseline models. For more detailed experimental findings, please refer to Table S.3 in the supplementary file.
@@ 4.5 Sensitivity Analysis of Hyperparameters $C_{1}$ and $C_{2}$
Figure 2: The effect of hyperparameter $(c_{1},c_{2})$ tuning on the accuracy (ACC) of some UCI and KEEL datasets on the performance of proposed GRVFL-MV model.
To gain a comprehensive understanding of how hyperparameters affect the generalization ability of the proposed GRVFL-MV model, we conducted a systematic exploration of the hyperparameter space by tuning the values of $c_{1}$ and $c_{2}$. This helps us to identify the optimal configuration that maximizes predictive accuracy and improves the model’s resilience to previously unseen data. Fig. 2. visually represents how the accuracy of the model behaves when the hyperparameters are tuned. The visual clearly shows that the proposed model is highly sensitive to the values of hyperparameters $c_{1}$ and $c_{2}$. In Fig. 2. (a), we see that optimal accuracy is achieved when $c_{1}=10^{5}$ and $c_{2}=10^{4}$ whereas, in Fig. 2. (c), optimal configuration is found at two distinct coordinates, $(10^{5},10^{5})$ and $(10,10^{5})$. Similar observations can be made for Fig. 2. (b) and Fig. 2. (d). Overall, these findings underscore the need for careful selection of hyperparameter values to achieve optimal model performance.